# ClinicalReTrial: A Self-Evolving AI Agent for Clinical Trial Protocol Optimization

### v1

@@ Abstract
Clinical trial failure remains a central bottleneck in drug development, where minor protocol design flaws can irreversibly compromise outcomes despite promising therapeutics. Although cutting-edge AI methods achieve strong performance in predicting trial success, they are inherently reactive for merely diagnosing risk without offering actionable remedies once failure is anticipated. To fill this gap, this paper proposes ClinicalReTrial, a self-evolving AI agent framework that addresses this gap by casting clinical trial reasoning as an iterative protocol redesign problem. Our method integrates failure diagnosis, safety-aware modification, and candidate evaluation in a closed-loop, reward-driven optimization framework. Serving the outcome prediction model as a simulation environment, ClinicalReTrial  enables low-cost evaluation of protocol modifications and provides dense reward signals for continuous self-improvement. To support efficient exploration, the framework maintains hierarchical memory that captures iteration-level feedback within trials and distills transferable redesign patterns across trials. Empirically, ClinicalReTrial  improves $83.3\%$ of trial protocols with a mean success probability gain of $5.7\%$, and retrospective case studies demonstrate strong alignment between the discovered redesign strategies and real-world clinical trial modifications.
ClinicalReTrial: A Self-Evolving AI Agent for Clinical Trial Protocol Optimization
Sixue Xing1, Xuanye Xia2, Kerui Wu3, Meng Jiang1, Jintai Chen4, Tianfan Fu5 1University of Notre Dame, South Bend, IN, USA 2Georgia Institute of Technology, Atlanta, GA, USA 3University of Massachusetts Amherst, Amherst, MA, USA 4Hong Kong University of Science and Technology (Guangzhou), Guangzhou, Guangdong, China 5Nanjing University, Nanjing, Jiangsu, China Correspondence: futianfan@gmail.com
@@ 1 Introduction
Clinical trials represent the most critical and expensive phase in drug discovery, with an estimated cost of \$2.6 billion (dimasi2016innovation) per approved drug, and low success rates of approximately 10-20% (yamaguchi2021approval). Serving as documented plan that specifies the study’s objectives, clinical trial protocols involve complex, interdependent design choices (getz2017trends), such as eligibility criteria, dosing strategies, and endpoint definitions, where small design flaws can propagate into irreversible failure. These challenges motivate the use of AI systems zhang2023harnessing that can reason over high-dimensional trial designs, leverage historical evidence, and systematically assess failure risks at scale.
@@
Recent advances in AI have enabled increasingly accurate prediction of clinical trial outcomes. For example, lo2019machine uses structured metadata to model success likelihood; fu2022hint; chen2024uncertainty; chen2025trialbench integrate heterogeneous data sources using architectures including graph neural networks and hierarchical attention mechanisms to achieve strong predictive performance; yue2024clinicalagent; liu2025autoct incorporate Large Language Models (LLMs) and external knowledge bases to enhance reasoning and explainability in trial outcome prediction.
@@
Despite their success, existing approaches are inherently reactive in nature: they operate on a fixed clinical trial protocol and produce a prediction or post-hoc explanation of trial success or failure. However, these methods do not address a more practically consequential problem: they are unable to respond to a determined trial failure due to the lack of actionable interventions. In real-world drug discovery, stakeholders require not only assessments of failure risk, but also actionable guidance on protocol redesign, including principled modifications or augmentations informed by the identified sources of risk.
@@
To bridge this gap, in this work, we propose ClinicalReTrial, a self-evolving AI agent framework that moves beyond static prediction toward actionable intervention via end-to-end trial protocol optimization, while continuously improves its redesign policies. The framework instantiates a coordinated multi-agent pipeline that performs failure diagnosis, protocol redesign, and candidate evaluation, with domain knowledge and safety awareness embedded at each decision stage. Beyond a single optimization run, we adopt the prediction model as a simulation environment to provide rewards for continuous self-improvement. Specifically, ClinicalReTrial  maintains local memory to accumulate iteration-level feedback and reward-attributed modification outcomes for within-trial adaptation, while a global memory distills transferable redesign patterns across trials to enable warm start initialization and exploration calibration. Through this hierarchical learning structure and reward-driven closed-loop optimization, ClinicalReTrial  systematically explores the protocol modification space and learns to identify high-impact interventions that improve clinical trial success probability.
@@
Experimentally, our prediction model demonstrated the strongest performance (PR-AUC >0.75>0.75), allowing it to serve as a reliable simulation environment for evaluation and agent optimization. In the trial redesign experiments, ClinicalReTrial successfully improve $89.3\%$ of trial protocols with mean probability gain $\Delta p=5.7\%$, achieved at negligible cost ($\mathdollar 0.12/trial$). We further conduct multiple real-world retrospective case studies. Impressively, the redesigns generated by ClinicalReTrial  exhibit strong strategic alignment with independently derived real-world trial modifications, highlighting the potential of self-evolving AI agents to support principled, clinically grounded trial redesign.
@@
Main contributions are listed as follows: (1) (to the best of our knowledge) We are the first to formulate clinical trial optimization as an AI-solvable and in silico-verifiable problem. (2) We propose a multi-agent pipeline with domain knowledge that decomposes clinical trial protocol optimization into analysis, augmentation, and evaluation. (3) We develop a simulation-driven clinical trial optimization framework with hierarchical memory utilization for continuous self-improvement.
@@ 2 Related Work
Early efforts employed classical machine learning (logistic regression (LR), random forests) on expert-curated features (gayvert2016clinical; lo2019machine), establishing feasibility but lacking multi-modal data integration. Deep learning approaches addressed this: fu2022hint proposed HINT, integrating drug molecules, ICD-10 codes, and eligibility criteria; chen2024uncertainty added uncertainty quantification; wang2024twin designed LLM-based patient-level digital twins; chen2025trialbench released a standardized TrialBench with multi-modal baselines. while maintaining competitive performance. Recent LLM approaches demonstrate medical reasoning (singhal2023medpalm), enhanced via retrieval-augmented generation (lewis2020rag) with databases like DrugBank (wishart2018drugbank), Hetionet (himmelstein2017hetionet), and domain-adapted encoders like BioBERT (lee2020biobert). Building on this, yue2024clinicalagent introduced ClinicalAgent, decomposing prediction into specialized sub-task agents with ReAct reasoning (yao2023react). liu2025autoct proposed AutoCT for autonomous feature engineering via Monte Carlo Tree Search (chi2024sela).
However, these methods function as discriminators mapping protocols to success probabilities without explaining why failures occur or how to modify protocols. First to formulate generative optimization, our multi-agent architecture leverages chain-of-thought (wei2022chain), and least-to-most prompting (zhou2023least) for hierarchical problem decomposition.

### v2

# ClinicalReTrial: Clinical Trial Redesign with Self-Evolving Agents

@@ Abstract
Clinical trials constitute a critical yet exceptionally challenging and costly stage of drug development (\$2.6B per drug), where protocols are encoded as complex natural language documents, motivating the use of AI systems beyond manual analysis. Existing AI methods accurately predict trial failure, but do not provide actionable remedies. To fill this gap, this paper proposes ClinicalReTrial, a multi-agent system that formulates clinical trial optimization as an iterative redesign problem on textural protocols. Our method integrates failure diagnosis, safety-aware modifications, and candidate evaluation in a closed-loop, reward-driven optimization framework. Serving the outcome prediction model as a simulation environment, ClinicalReTrial enables low-cost evaluation and dense reward signals for continuous self-improvement. We further propose a hierarchical memory that captures iteration-level feedback within trials and distills transferable redesign patterns across trials. Empirically, ClinicalReTrial improves $83.3\%$ of trial protocols with a mean success probability gain of $5.7\%$ with negligible cost (\$0.12 per trial). Retrospective case studies demonstrate alignment between the discovered redesign strategies and real-world clinical trial modifications. The code is anonymously available at: https://github.com/xingsixue123/ClinicalFailureReasonReTrial.
ClinicalReTrial: Clinical Trial Redesign with Self-Evolving Agents
Sixue Xing1, Kerui Wu2, Xuanye Xia3, Meng Jiang1, Jintai Chen4, Tianfan Fu5 1University of Notre Dame, South Bend, IN, USA 2University of Massachusetts Amherst, Amherst, MA, USA 3Georgia Institute of Technology, Atlanta, GA, USA 4Hong Kong University of Science and Technology (Guangzhou), Guangzhou, Guangdong, China 5Nanjing University, Nanjing, Jiangsu, China Correspondence: futianfan@gmail.com
@@ 1 Introduction
Clinical trials represent the most critical and expensive phase in drug discovery, with an estimated cost of \$2.6 billion (DiMasi et al., 2016) per approved drug, and low success rates of approximately 10-20% (Yamaguchi et al., 2021). Serving as documented plan that specifies the study’s objectives, clinical trial protocols involve complex, interdependent design choices (Getz and Campo, 2017) expressed in natural language documents, such as eligibility criteria, dosing strategies, and endpoint definitions, where small design flaws can propagate into irreversible failure. These challenges motivate the use of AI systems Zhang et al. (2023) that can reason over high-dimensional trial designs, leverage historical evidence, and systematically assess failure risks at scale.
@@
Recent advances in AI have enabled increasingly accurate prediction of clinical trial outcomes. For example, Lo et al. (2019) uses structured metadata to model success likelihood; Fu et al. (2022); Chen et al. (2024c, 2025) integrate heterogeneous data sources using architectures including graph neural networks and hierarchical attention mechanisms to predict trial approval rate; Yue et al. (2024); Liu et al. (2025) incorporate Large Language Models (LLMs) and external knowledge bases to enhance reasoning and explainability in trial outcome prediction through natural advanced language understandings of complex medical texts.
@@
Despite their success, existing approaches are inherently reactive in nature: they operate on a fixed clinical trial protocol and produce a prediction or post-hoc explanation of trial success or failure. However, these methods do not address a more practically consequential problem: they are unable to respond to a determined trial failure due to the lack of actionable interventions. In real-world drug discovery, stakeholders require not only assessments of failure risk, but also actionable guidance on protocol redesign, including principled modifications or augmentations informed by the identified sources of risk (Dagenais et al., 2022; Baumfeld Andre et al., 2020).
Figure 1: ClinicalReTrial Agent architecture. The system operates through iterative refinement: agents analyze failures, generate modifications, and receive rewards from the simulation environment. Historical explorations are extracted into structured knowledge that guides subsequent iterations, enabling progressive improvement.
@@
To bridge this gap, in this work, as shown in Figure˜1, we propose ClinicalReTrial, a self-evolving AI agent that moves beyond static prediction toward actionable intervention via end-to-end trial protocol optimization, while continuously improving its redesign policies. The inherently language-rich nature of clinical trial protocols makes this an ideal domain for LLM-based optimization. Our framework instantiates a coordinated multi-agent pipeline that performs failure diagnosis, protocol redesign, and candidate evaluation, with domain knowledge and safety awareness embedded at each decision stage. Beyond a single optimization run, we adopt the prediction model as a simulation environment to provide rewards for continuous self-improvement. Specifically, ClinicalReTrial maintains local memory to accumulate iteration-level feedback and reward attributed modification outcomes for within-trial adaptation, while a global memory distills transferable redesign patterns across trials to enable warm start initialization and exploration calibration. Through this hierarchical learning structure and reward-driven closed-loop optimization, ClinicalReTrial systematically explores the protocol modification space and learns to identify high-impact interventions that improve clinical trial success probability.
@@
Experimentally, our prediction model demonstrated the strongest performance (PR-AUC>0.75>0.75), allowing it to serve as a reliable simulation environment for evaluation and agent optimization. In the trial redesign experiments, ClinicalReTrial successfully improved $83.3\%$ of trial protocols with mean probability gain $\Delta p=5.7\%$, achieved at negligible cost ($\mathdollar 0.12/$trial). We further conduct multiple real-world retrospective case studies. Impressively, the redesigns generated by ClinicalReTrial exhibit strategic alignment with independently derived real-world trial modifications, highlighting the potential of self-evolving AI agents to support principled trial redesign.
Main contributions.
@@
are listed as follows: (1) (to the best of our knowledge) We are the first to formulate clinical trial optimization as an AI-solvable and in silico-verifiable problem. (2) We propose a multi-agent pipeline with domain knowledge that decomposes clinical trial protocol optimization into diagnosis, modification, and evaluation. (3) We develop a simulation-driven clinical trial optimization framework with In-Context Learning and multi-level memory for continuous self-improvement through reward attributed prompt optimization, dynamic redesign pool curation, and cross-trial knowledge distillation.
@@ 2 Related Work
Early efforts employed classical machine learning (logistic regression (LR), random forests) on expert-curated features (Gayvert et al., 2016; Lo et al., 2019), establishing feasibility but lacking multi-modal data integration. Deep learning approaches addressed this: Fu et al. (2022) proposed HINT, integrating drug molecules, ICD-10 codes, and eligibility criteria; Chen et al. (2024c) added uncertainty quantification and interpretability; Wang et al. (2024) designed LLM-based patient-level digital twins; Chen et al. (2025) released a standardized TrialBench with multi-modal baselines while maintaining competitive performance. Recent LLM approaches demonstrate medical reasoning (Singhal et al., 2023), enhanced via retrieval-augmented generation (Lewis et al., 2020) with databases like DrugBank (Wishart et al., 2018), Hetionet (Himmelstein et al., 2017), and domain-adapted encoders like BioBERT (Lee et al., 2020). Building on this, Yue et al. (2024) introduced ClinicalAgent, decomposing prediction into specialized sub-task agents with ReAct reasoning (Yao et al., 2023). Liu et al. (2025) proposed AutoCT for autonomous feature engineering via Monte Carlo Tree Search (Chi et al., 2024). However, these methods are essentially predictive models without explaining why failures occur or how to modify protocols. In contrast, our multi-agent architecture leverages chain-of-thought (Wei et al., 2022) and least-to-most prompting (Zhou et al., 2023) to dynamically diagnose failure causes and refine protocols.